OpenStudy (anonymous):

What is the sum of a 14–term arithmetic sequence where the last term is 30 and the common difference is –5? 875 422.5 812.5 455

OpenStudy (anonymous):

a[n] = a[1] + (n-1)d = a[1] - 5(n-1) a[14] = a[1] - 5(14-1) = a[1] - 65 30 = a[1] - 65 a[1] = 95 sum of first 14 terms = 14(a[1] + a[14])/2 = 14(95 + 30)/2 = 875